OpenStudy (anonymous):

consider f(x) = 2x^2-3x-2 use factoring to solve for the zeros of f

OpenStudy (anonymous):

(2x+1)(x-2) x=-(1/2) and 2

OpenStudy (anonymous):

is that all?

OpenStudy (anonymous):

that should be all i think haha

OpenStudy (anonymous):

how did you do this problem just so i know how to do it?

OpenStudy (anonymous):

so, there are two ways 1) using the quadratic formula 2) (this one is probably what you Should be doing at this point seems like) is to use trial and error to see *what* multiplied by *what* gives you x^2-3x-4 if you multiply (2x+1)(x-2), you get x^2-3x-4

OpenStudy (anonymous):

so the ultimate goal of this problem is to find the zeros you have to make it into this equation x^2-3x-4=0 then (2x+1)(x-2)=0 then what "x" makes a zero? -(1/2) and 2

OpenStudy (anonymous):

oh ok i see

OpenStudy (anonymous):

i saw that part but what did you do to get (2x+1) and (x-2)

OpenStudy (anonymous):

sorry for the late reply.. you just have to try and see what works with practice, you can get it more quickly/see commonalities in problems

OpenStudy (anonymous):

i mean what did you do to get it? that is probably the same way i would do it.

OpenStudy (anonymous):

erm.. well, i guess when i see something like 2x^2-3x-2 i first put (2x )(x ) then i look at the last number and see that it's got to be -2*1 or 2*-1 then it's just matter of which works (2x+2)(x-1) or (2x-1)(x+2)

OpenStudy (anonymous):

so its what multiples to -2 and adds to -3?

OpenStudy (anonymous):

o oops, in hast, i put the wrong factors in my last post, but you understand i'm not sure what you're asking but i guess yes the middle term is the result of the addition of the 2nd terms in the factor (if there's no coefficient in the first term) so let's say (x+1)(x-2) the middle term would be 1x (as -2+1 is 1) actually.. i might be making this more confusing do you understand so far?

OpenStudy (anonymous):

im beginning to understand it

OpenStudy (anonymous):

so its not 2x+1

OpenStudy (anonymous):

?

OpenStudy (anonymous):

you mean "so one of the factors isn't (2x+1)"? it is if you multiply through.. (2x+1)(x-2) = 2x*x + 2x*-2 + 1*x + 1*-2 =2x^2 - 4x + x - 2 = 2x^2 - 3x - 2

OpenStudy (anonymous):

if it's too confusing now...you might just be best advised to learn it in class where it may be better explained